OpenStudy (anonymous):

Write the expression as a single natural logarithm. 2 ln x-4 ln c

OpenStudy (anonymous):

first use \(\large nlog_bM=log_bM^n \) then use \(\large log_bM-log_bN=log_b{\frac{M}{N}} \)

OpenStudy (anonymous):

That is extremely confusing haha is there any way you could help me make sense of it?

OpenStudy (zepp):

Is it \(2\ln(x-4)\ln(c)\)?

OpenStudy (zepp):

Or \(2\ln(x)-4\ln(c)\)?

OpenStudy (anonymous):

The choices are ln X^2/c^4 ln c^4/x^2 ln (x^2 +c^4) and ln (x^2-c^4)